OpenStudy (help_people):

Which of the following functions has the largest value when x = 4? e(x) = x2 + 6x + 21 m(x) = 8x v(x) = 31x e(x) v(x) m(x) All the functions are equal at x = 4.

OpenStudy (help_people):

c?

OpenStudy (help_people):

@triciaal

Mehek (mehek14):

the only exponent is \(x^2\) right?

OpenStudy (help_people):

yes :)

OpenStudy (help_people):

8x is 8^x also 31x is 31(x)

Mehek (mehek14):

\(x^2\\8^x\\31x\) like that?

OpenStudy (anonymous):

plug 4 in each of the functions and see which is the largest value

OpenStudy (help_people):

yes

OpenStudy (help_people):

i did i got c but I'm making sure

Mehek (mehek14):

i'm asking if it's like this \(x^2\\8^x\\31x\)

OpenStudy (help_people):

yes it is @Mehek14 i already said that :)

OpenStudy (anonymous):

if this is what u mean e(x) = x^2 + 6x + 21 m(x) = 8^x v(x) = 31x then when we plug in 4 we get e(x)= 61 m(x)= 4096 v(x)= 124

OpenStudy (anonymous):

meaning m(x) is the highest value

Mehek (mehek14):

C is correct

OpenStudy (help_people):

ok i thought so thank you everybody ! :)

OpenStudy (anonymous):

you're welcome :D

OpenStudy (triciaal):

yes C